1. Find the greatest common divisor of the following polynomials over F, the field of rational numbers: (a) x 3 - 6x 2 + x + 4 and x 5 - 6x + 1. (b) x 2 + 1 and x6 + x 3 + x + 1. 


1
Expert's answer
2022-03-02T12:42:25-0500

(a)

"f(x)=x^3-6x^2+x+4\\
g(x)=x^5-6x+1\\
g:f\\
\begin{matrix}
 x^5-6x+1 &|x^3-6x^2+x+4 \\
 x^5-6x^4+x^3+4x^2 & x^2+6x+35\\
---------\\
6x^4-x^3-4x^2-6x+1\\
6x^4-36x^3+6x^2+24x\\
-----------\\
35x^3-10x^2-30x+1\\
35x^3-210x^2+35x+140\\
----------\\
200x^2-65x-139
\end{matrix}"

"r=200x^2-65x-139\\
f:r"

"\begin{matrix}
 x^3-6x^2+x+4 & |200x^2-65x-139 \\
 ------- & x-1135\\
200x^3-1200x^2+200x+800\\
200x^3-65x^2-139x\\
--------\\
-1135x^2+339x+800\\
---------\\
-22700x^2+67800x+160000\\
-22700x^2+73775x+157765\\
--------\\
-5975x+2235\\
------\\
-1195x+447
\end{matrix}"

"r_1=-1195x+447\\
r:r_1\\
\begin{matrix}
 200x^2-65x-139 &|-1195x+447 \\
 -------& 200x+2345\\
-23900x^2+77675x+166105\\
-23900x^2+89400x\\
-------\\
-11725x+166105\\
--------\\
2345x+33221\\
------\\
-2802275x-39699095\\
-2802275x-1048215\\
-----\\
-38650880\neq0
\end{matrix}\\
gcd(f,g)=1"

(b)

"f(x)=x^6+x^3+x+1\\
g(x)=x^2+1\\
f:g\\
\begin{matrix}
 x^6+x^3+x+1& |x^2+1 \\
 x^6+x^4 & x^4-x^2+x+1\\
-----\\
-x^4+x^3+x+1\\
-x^4-x^2\\
-----\\
x^3+x^2+x+1\\
x^3+x\\
-----\\
x^2+1\\
x^2+1\\
-----\\0
\end{matrix}\\
gcd(f,g)=x^2+1"